OpenStudy (anonymous):

Bernard is asked to create a swimming pool and deck design for a new series of model homes currently being designed. Each home will have varying pool sizes depending on the lot size; however, there will be some similarities, so polynomial expressions will be used. Your task will be to help Bernard. To fit the dimensions of the various land spaces, he realizes that the length of the pool has to be three times the size of the width plus 3, and the depth has to be 7 less than twice the length. @ganeshie8 @Abhisar @Luigi0210 @paki

OpenStudy (anonymous):

a. Create a variable to represent the width of the pool. Determine the expressions to represent length and the depth in terms of the width variable. b. Determine the polynomial expression to represent the area of land space that the pool will cover. c. Using the area and the depth expressions, determine the polynomial expression that is used to represent the volume. d. Classify the volume polynomial by degree and number of terms.

OpenStudy (anonymous):

One second let me look at it

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Wow this is a lot of explaining to do...

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

Create a variable to represent the width of the pool. Determine the expressions to represent length and the depth in terms of the width variable. Length = 3w + 3 Depth = 2l – 7 Width = 1/3 l-1

OpenStudy (anonymous):

so thats a?

OpenStudy (anonymous):

One second let me do it again

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

a) Length = 3w +3 = 21 Depth = 21 - 7 Width = 1/31 - 1

OpenStudy (anonymous):

b) Surface Area of Rectangular Prism = 2(wh+lw+lh) 2(((1/3 l-1)∙(2l-7) )+((3w+3)∙(1/3 l-1) )+

OpenStudy (anonymous):

Dang Thanks.

OpenStudy (anonymous):

You want help with the others?

OpenStudy (anonymous):

Yes and is that the answer or B?

OpenStudy (anonymous):

Here is the answer for b)

OpenStudy (anonymous):

Surface Area of Rectangular Prism = 2(wh+lw+lh) 2(1/3 l-1) ∙ (2l-7) ) + (3w+3) ∙ (1/3 l-1) )

OpenStudy (anonymous):

Huhh.. why is it like this...

OpenStudy (anonymous):

refresh the page

OpenStudy (anonymous):

Surface Area of Rectangular Prism = 2(wh+lw+lh) 2(1/31- 1) x (21 - 7) + (3w + 3) x (1/31- 1)

OpenStudy (anonymous):

There we go

OpenStudy (anonymous):

Okay now how about C and D

OpenStudy (anonymous):

The answer for c)

OpenStudy (anonymous):

Btw thanks so much!!

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

The volume of a rectangular prism is length times width times height. (3w + 3) x (21 - 7) x (1/31 - 1)

OpenStudy (anonymous):

That's the answer for c)

OpenStudy (anonymous):

For d)

OpenStudy (anonymous):

d) I simplified the volume and it came out to. 〖21l〗^2 w+13lw+〖2l〗^2+21w-13l+21 This is a third degree polynomial.

OpenStudy (anonymous):

nope let me do it again

OpenStudy (anonymous):

if you refresh the page it gets rid of the question marks.

OpenStudy (anonymous):

Ohh thanks

OpenStudy (anonymous):

But I think that's all

OpenStudy (anonymous):

okay thanks!!

OpenStudy (anonymous):

Ur welcome... That was a lot of work

OpenStudy (anonymous):

what are them things that is by the 211 or somthing

OpenStudy (anonymous):

o just well u can do it like this instead (

OpenStudy (anonymous):

I mean ( or )

OpenStudy (anonymous):

Is that good now?

OpenStudy (anonymous):

Yes but we got a problem there is more ill tag you in a new thread